# Hierarchical Relation-augmented Representation Generalization for Few-shot Action Recognition

### v1

@@ 3.3. Inter-video Semantic Correlation
Different from exploiting inter-frame interactions in each video, capturing inter-video relationships can strengthen task-specific semantic patterns for each video. To achieve this goal, we design Inter-video Semantic Correlation (ISC) module to conduct Fine-grained Inter-video Interaction and further learn intra- and inter-class correlation by Adaptive Interaction Mask (Fig. 3(a)) for support-query and support-support videos (Fig. 2(b)). Via Adaptive Interaction Mask, support and query videos aggregate features from different types of videos (i.e., support or query) based on their respective characteristics. Compared with previous works (Wang et al., 2022; Xing et al., 2023c), which only consider utilize global video features to perform interactions among support and query videos, we attempt to make use of temporal cues in the video to strengthen sufficient temporal pattern understanding among videos.
@@
Fine-grained Inter-video Interaction. Based on the above ananlysis, how to perform fine-grained inter-video interactions is crucial for FSAR. A straightforward approach is to perform dense attention along temporal and sample dimensions simultaneously. Thus, given a task with $NK$ support videos and $L$ query videos, the computational complexity of such dense sample attention is $O((NK+L)^{2}T^{2})$, causing high computational costs. Besides, due to the difference between the temporal dimension and the sample dimension, it is not reasonable to treat them equivalently.
@@
Thus, we decouple dense sample attention into inter-frame interaction and our designed Inter-video Semantic Correlation (ISC). The inter-frame interaction is achieved by previous Inter-frame Temporal Modeling, obtaining support and query temporal features, i.e., $\bm{F}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$ and $\bm{F}_{\mathrm{q}}\in{\mathbb{R}}^{L\times T\times C}$. Next, ISC concatenates support and query features, and reshapes these features in one task to ${\mathbb{R}}^{T\times(NK+L)\times C}$. Afterwards, multi-head self-attention along with sample attention matrix $\bm{A}\in{\mathbb{R}}^{(NK+L)\times(NK+L)}$ is adopted to conduct cross-video interactions by attending all frame features at the same temporal location, resulting in task-specific temporal futures for support and query videos:
(1) $\tilde{\bm{F}}=[\tilde{\bm{F}}_{\mathrm{s}},\tilde{\bm{F}}_{\mathrm{q}}]=\texttt{SelfAttn}(\texttt{Re}(\texttt{Cat}(\bm{F}_{\mathrm{s}},\bm{F}_{\mathrm{q}}))),\vspace{-2pt}$
@@
Adaptive Interaction Mask. Due to the difference between support and query videos, we propose Adaptive Interaction Mask $\bm{J}\in{\mathbb{R}}^{(NK+L)\times(NK+L)}$ for support-support and query-support interactions, as shown in Fig. 3(a). Specifically in each task, support and query videos only aggregate semantic information from support videos while discarding interactions with query videos by element-wise multiplication of $\bm{A}$ and $\bm{J}$. By transferring knowledge between support-support and query-support videos, support features could learn intra- and inter-class correlation, and query features capture task-specific temporal cues. Besides, as shown in Fig. 3(b)(c)(d), we try other masked interaction strategies for inter-video interactions, including support-support interaction, query-support interaction and full interaction. Specifically, support-support interaction denotes ISC only considers the interactions among support videos, while query-support interaction denotes query videos only aggregate semantic information from support videos in ISC. full interaction means ISC simultaneously considers the interactions among query and support videos, but ignores the distinct characteristics of the query and support videos. The experiments about different strategies are provided in Table 4(a).
Figure 3. Different masked interaction strategies for Inter-video Semantic Correlation.
3.4. Inter-task Knowledge Transfer
@@
Prototype Construction. The core idea is to incorporate frame-level features into compact temporal prototypes, so as to filter out redundant information and obtain key temporal cues. We first construct $M$ learnable prototypes $\bm{P}\in{\mathbb{R}}^{M\times C}$ to summarize video representations, where we empirically set $M$ to 3 (ablation study in Table 4(c)). Given a task consisting of support and query videos, these prototypes adaptively aggregate action dynamic information from support features $\tilde{\bm{F}}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$:
(2) $\hat{\bm{P}}=\mathrm{Softmax}(\bm{P}\bm{K}_{\mathrm{t}}^{\top})\bm{V}_{\mathrm{t}}+\bm{P},$
@@
Knowledge Retrieval. Inspired by (Wu et al., 2018; Zhou et al., 2022), we setup a temporal knowledge bank $\bm{M}=[\bm{m}_{1},\bm{m}_{2},...,\bm{m}_{G}]\in{\mathbb{R}}^{G\times C}$ to store task-shared temporal patterns, where $G$ is the number of memory representations and $G$ is empirically set to $50$ (ablation study in Fig. 4). After obtaining original temporal prototypes $\hat{\bm{P}}$, we perform knowledge retrieval via only aggregating semantic-related memory from the bank:
(3) $\bm{P}^{\prime}=\sum_{g=1}^{G}\mathrm{topk}(\hat{\bm{P}},\bm{m}_{g})\bm{m}_{g},$
where $\mathrm{topk}$ is the function that retrieves $O$ most similar prototypes from the bank according to the semantic correspondences. We take a hyper-parameter $\kappa\!=\!O/G$ to determine the number of retrieved memory representations. $\bm{P}^{\prime}$ is retrieved historical knowledge related to current task, which could be used to further enhance task-specific video representations.
@@
(4) $\small\bar{\bm{F}}_{\mathrm{s}}=\texttt{CrossAttn}(\tilde{\bm{F}}_{\mathrm{s}},\bar{\bm{P}}),$
@@
(5) $\bm{m}_{t}=\texttt{SelectMax}\left(\frac{\bm{p}_{i}\cdot\bm{m}_{j}}{||\bm{p}_{i}||||\bm{m}_{j}||}\right),\bm{m}_{j}\in\bm{M},$
(6) $\bm{m}_{t}\leftarrow\mu\bm{m}_{t}+(1-\mu)\bm{p}_{i},$
where SelectMax denote the function that selects most similar memory representation from $\bm{M}$ as $\bm{m}_{t}$, and $\mu\!\in\![0,1]$ is the momentum for memory evolution. Note that during the testing phase, the temporal knowledge bank remains unchanged and does not store any information from testing data, avoiding information leakage.
@@
3.5. Training and Testing with HR2G-shot
Training Objective. After jointly exploring three types of relation modeling, we obtain support features $\bar{\bm{F}}_{\mathrm{s}}$ and query features $\bar{\bm{F}}_{\mathrm{q}}$. Given $i$-th support frame-level features $\bar{\bm{f}}_{\mathrm{s}}^{i}\in{\mathbb{R}}^{T\times C}$ in $\bar{\bm{F}}_{\mathrm{s}}$ and $j$-th query frame-level features $\bar{\bm{f}}_{\mathrm{q}}^{j}\in{\mathbb{R}}^{T\times C}$ in $\bar{\bm{F}}_{\mathrm{q}}$, like in previous methods (Cao et al., 2020; Wang et al., 2024b), we adopt the “temporal alignment metric to obtain query-support distances:
(7) ${\mathcal{D}}=\mathtt{Metric}(\bar{\bm{f}}_{\mathrm{q}}^{j},\bar{\bm{f}}_{\mathrm{s}}^{i}),$
where $\mathtt{Metric}$ denotes the OTAM (Cao et al., 2020) metric by default. Then we can use the output support-query distances as logits to compute cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$ over the ground-truth labels.
@@ 3.5. Training and Testing with HR2G-shot
Table 1. Quantitative results on SSv2-small (Goyal et al., 2017), SSv2-full (Goyal et al., 2017), HMDB51 (Kuehne et al., 2011), UCF101 (Soomro et al., 2012), and Kinetics (Carreira and Zisserman, 2017) (see §4.3). “INet-RN50” denotes ResNet-50 pre-trained on ImageNet. The best results are highlighted.
  SSv2-small SSv2-full HMDB51 UCF101 Kinetics
Method Reference Pre-training $1$-shot $5$-shot $1$-shot $5$-shot $1$-shot $5$-shot $1$-shot $5$-shot $1$-shot $5$-shot
CMN (Zhu and Yang, 2018) ECCV’18 INet-RN50 $34.4$ $43.8$ $36.2$ $48.9$ $60.5$ $78.9$
OTAM (Cao et al., 2020) CVPR’20 INet-RN50 - - $42.8$ $52.3$ $73.0$ $85.8$
AmeFuNet (Fu et al., 2020) MM’20 INet-RN50 - - - - $60.2$ $75.5$ $85.1$ $95.5$ $74.1$ $86.8$
TRX (Perrett et al., 2021) CVPR’21 INet-RN50 - $59.1$ $64.6$ $75.6$ $96.1$ $63.6$ $85.9$
TA2(Li et al., 2022) AAAI’22 INet-RN50 - - $47.6$ $61.0$ $59.7$ $73.9$ $81.9$ $95.1$ $72.8$ $85.8$
MTFAN (Wu et al., 2022) CVPR’22 INet-RN50 - - $45.7$ $60.4$ $59.0$ $74.6$ $84.8$ $95.1$ $74.6$ $87.4$
HyRSM (Wang et al., 2022) CVPR’22 INet-RN50 $40.6$ $56.1$ $54.3$ $69.0$ $60.3$ $76.0$ $83.9$ $94.7$ $73.7$ $86.1$
STRM (Thatipelli et al., 2022) CVPR’22 INet-RN50 $37.1$ $55.3$ $43.1$ $68.1$ $52.3$ $77.3$ $80.5$ $96.9$ $62.9$ $86.7$
HCL (Zheng et al., 2022) ECCV’22 INet-RN50 $38.9$ $55.4$ $47.3$ $64.9$ $59.1$ $76.3$ $82.6$ $94.5$ $73.7$ $85.8$
MoLo (Wang et al., 2023a) CVPR’23 INet-RN50 $41.9$ $56.2$ $55.0$ $69.6$ $60.8$ $77.4$ $86.0$ $95.5$ $74.0$ $85.6$
GgHM (Xing et al., 2023c) ICCV’23 INet-RN50 - - $54.5$ $69.2$ $85.2$ $96.3$ $61.2$ $76.9$ $74.9$ $87.4$
CLIP-FSAR (Wang et al., 2024b) IJCV’24 CLIP-RN50 $52.0$ $55.8$ $58.1$ $62.8$ $69.2$ $80.3$ $91.3$ $97.0$ $87.6$ $91.9$
CLIP-Freeze (Radford et al., 2021) ICML’21 CLIP-ViT-B $29.5$ $42.5$ $30.0$ $42.4$ $58.2$ $77.0$ $89.7$ $95.7$ $78.9$ $91.9$
CapFSAR (Wang et al., 2023b) Arxiv’23 CLIP-ViT-B $45.9$ $59.9$ $51.9$ $68.2$ $65.2$ $78.6$ $93.3$ $97.8$ $84.9$ $93.1$
MVP-shot (Qu et al., 2024) Arxiv’24 CLIP-ViT-B $55.4$ $62.0$ $77.0$ $88.1$ $96.8$ $99.0$ $91.0$ $95.1$
MA-FSAR (Xing et al., 2024) Arxiv’24 CLIP-ViT-B $59.1$ $64.5$ $63.3$ $72.3$ $83.4$ $87.9$ $97.2$ $99.2$ $\mathbf{95.7}$ $96.0$
CLIP-FSAR (Wang et al., 2024b) IJCV’24 CLIP-ViT-B $54.6$ $61.8$ $62.1$ $72.1$ $77.1$ $87.7$ $97.0$ $99.1$ $94.8$ $95.4$
HR2G-shot (ours) - CLIP-ViT-B $\mathbf{60.2}$ $\mathbf{66.0}$ $\mathbf{65.4}$ $\mathbf{74.8}$ $\mathbf{85.6}$ $\mathbf{88.6}$ $\mathbf{98.0}$ $\mathbf{99.3}$ 95.2 $\mathbf{96.4}$

### v3

@@ 3.3 Inter-video Semantic Correlation
Different from exploiting inter-frame interactions in each video, capturing inter-video relationships can strengthen task-specific semantic patterns for each video. To achieve this goal, we design Inter-video Semantic Correlation (ISC) module to conduct Fine-grained Inter-video Interaction and further learn intra- and inter-class correlation by Adaptive Interaction Mask (Fig. 3(a)) for support-query and support-support videos (Fig. 2(b)). Via Adaptive Interaction Mask, support and query videos aggregate features from different types of videos (i.e., support or query) based on their respective characteristics. Compared with previous works [13, 15], which only consider utilizing global video features to perform interactions among support and query videos, we attempt to make use of temporal cues in the video to strengthen sufficient temporal pattern understanding among videos.
Figure 3: Different masked interaction strategies for Inter-video Semantic Correlation.
@@
Fine-grained Inter-video Interaction. Based on the above analysis, how to perform fine-grained inter-video interactions is crucial for FSAR. A straightforward approach is to perform dense attention along temporal and sample dimensions simultaneously. Thus, given a task with $NK$ support videos and $L$ query videos, the computational complexity of such dense sample attention is $O((NK+L)^{2}T^{2})$, causing high computational costs. Besides, due to the difference between the temporal dimension and the sample dimension, it is not reasonable to treat them equivalently.
@@
Thus, we decouple dense sample attention into inter-frame interaction and our designed Inter-video Semantic Correlation (ISC). The inter-frame interaction is achieved by previous Inter-frame Temporal Modeling, obtaining support and query temporal features, i.e., $\bm{F}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$ and $\bm{F}_{\mathrm{q}}\in{\mathbb{R}}^{L\times T\times C}$. Next, ISC concatenates support and query features, and reshapes these features in one task to ${\mathbb{R}}^{T\times(NK+L)\times C}$. Afterwards, multi-head self-attention along with sample attention matrix $\bm{A}\in{\mathbb{R}}^{(NK+L)\times(NK+L)}$ is adopted to conduct cross-video interactions by attending all frame features at the same temporal location, resulting in task-specific temporal features for support and query videos:
$\tilde{\bm{F}}=[\tilde{\bm{F}}_{\mathrm{s}},\tilde{\bm{F}}_{\mathrm{q}}]=\texttt{SelfAttn}(\texttt{Re}(\texttt{Cat}(\bm{F}_{\mathrm{s}},\bm{F}_{\mathrm{q}}))),$ (1)
@@
Adaptive Interaction Mask. Due to the difference between support and query videos, we propose Adaptive Interaction Mask $\bm{J}\in{\mathbb{R}}^{(NK+L)\times(NK+L)}$ for support-support and query-support interactions, as shown in Fig. 3(a). Specifically in each task, support and query videos only aggregate semantic information from support videos while discarding interactions with query videos by element-wise multiplication of $\bm{A}$ and $\bm{J}$. By transferring knowledge between support-support and query-support videos, support features could learn intra- and inter-class correlation, and query features capture task-specific temporal cues. Besides, as shown in Fig. 3(b)(c)(d), we try other masked interaction strategies for inter-video interactions, including support-support interaction, query-support interaction, and full interaction. Specifically, support-support interaction denotes that ISC only considers the interactions among support videos, while query-support interaction denotes query videos only aggregate semantic information from support videos in ISC. full interaction means ISC simultaneously considers the interactions among query and support videos, but ignores the distinct characteristics of the query and support videos. The experiments about different strategies are provided in Table 6.
Algorithm S1 Pseudo-code for Inter-video Semantic Correlation and Inter-task Knowledge Transfer in a PyTorch-like style.
# Inter-video Semantic Correlation
def InterVideoCorrelation(support_feats, query_feats):
# Fine-grained Inter-video interaction
# support_feats: [N*K, T, D], query_feats: [Q, T, D]
support_query = Reshape(Cat(support_feats, query_feats)) # [T, N*K+Q, D]
# Aggregating support information to refine query representation
# adaptive_mask: [N*K+Q, N*K+Q]
support_query = SelfAttn(support_query, adaptive_mask)
support_feats, query_feats = Reshape(support_query)
return support_feats, query_feats
# Inter-task Knowledge Transfer
def InterTaskTransfer(support_feats, prototypes):
# Incorporate frame-level features into compact prototypes
# support_feats: [N*K, T, D], prototypes: [M, D]
prototypes = CrossAttn(prototypes, support_feats)
# Aggregate semantic-related memory into prototypes from the temporal bank
# temporal_bank: [G, D]
prototypes = KnowledgeRetrieval(prototypes, temporal_bank)
# Aggregate temporal knowledge into support features via cross attention
support_feats = CrossAttn(support_feats, prototypes)
# Update the temporal bank using the prototypes
temporal_bank = UpdatingBank(temporal_bank, prototypes)
return support_feats, temporal_bank
3.4 Inter-task Knowledge Transfer
@@
Prototype Construction. The core idea is to incorporate frame-level features into compact temporal prototypes, so as to filter out redundant information and obtain key temporal cues. We first construct $M$ learnable prototypes $\bm{P}\in{\mathbb{R}}^{M\times C}$ to summarize video representations, where we empirically set $M$ to 3 (ablation study in Table 8). Given a task consisting of support and query videos, these prototypes adaptively aggregate action dynamic information from support features $\tilde{\bm{F}}_{\mathrm{s}}\in{\mathbb{R}}^{NK\times T\times C}$:
$\hat{\bm{P}}=\mathrm{Softmax}(\bm{P}\bm{K}_{\mathrm{t}}^{\top})\bm{V}_{\mathrm{t}}+\bm{P},$ (2)
@@
Knowledge Retrieval. Inspired by [62, 63], we set up a temporal knowledge bank $\bm{M}=[\bm{m}_{1},\bm{m}_{2},...,\bm{m}_{G}]\in{\mathbb{R}}^{G\times C}$ to store task-shared temporal patterns, where $G$ is the number of memory representations and $G$ is empirically set to $50$ (ablation study in Fig. 4). After obtaining original temporal prototypes $\hat{\bm{P}}$, we perform knowledge retrieval via only aggregating semantic-related memory from the bank:
$\bm{P}^{\prime}=\sum_{g=1}^{G}\mathrm{topk}(\hat{\bm{P}},\bm{m}_{g})\bm{m}_{g},$ (3)
where $\mathrm{topk}$ is the function that retrieves $O$ most similar prototypes from the bank according to the semantic correspondences. We take a hyper-parameter $\kappa\!=\!O/G$ to determine the number of retrieved memory representations. $\bm{P}^{\prime}$ is retrieved historical knowledge related to the current task, which could be used to further enhance task-specific video representations.
@@
$\small\bar{\bm{F}}_{\mathrm{s}}=\texttt{CrossAttn}(\tilde{\bm{F}}_{\mathrm{s}},\bar{\bm{P}}),$ (4)
@@
[…]
$\bm{m}_{t}\leftarrow\mu\bm{m}_{t}+(1-\mu)\bm{p}_{i},$ (6)
where SelectMax denotes the function that selects most similar memory representation from $\bm{M}$ as $\bm{m}_{t}$, and $\mu\!\in\![0,1]$ is the momentum for memory evolution. Note that during the testing phase, the temporal knowledge bank remains unchanged and does not store any information from testing data, avoiding information leakage.
@@
Algorithm S1 presents the pseudo-code for Inter-video Semantic Correlation (§3.3) and Inter-task Knowledge Transfer (§3.4) in a PyTorch-like style.
3.5 Training and Testing with HR2G-shot
Training Objective. After jointly exploring three types of relation modeling, we obtain support features $\bar{\bm{F}}_{\mathrm{s}}$ and query features $\bar{\bm{F}}_{\mathrm{q}}$. Given $i$-th support frame-level features $\bar{\bm{f}}_{\mathrm{s}}^{i}\in{\mathbb{R}}^{T\times C}$ in $\bar{\bm{F}}_{\mathrm{s}}$ and $j$-th query frame-level features $\bar{\bm{f}}_{\mathrm{q}}^{j}\in{\mathbb{R}}^{T\times C}$ in $\bar{\bm{F}}_{\mathrm{q}}$, like in previous methods [10, 61], we adopt the temporal alignment metric to obtain query-support distances:
${\mathcal{D}}=\mathtt{Metric}(\bar{\bm{f}}_{\mathrm{q}}^{j},\bar{\bm{f}}_{\mathrm{s}}^{i}),$ (7)
where $\mathtt{Metric}$ denotes the OTAM [10] metric by default. Then we can use the output support-query distances as logits to compute cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$ over the ground-truth labels.
@@ 3.5 Training and Testing with HR2G-shot
Table 1: Quantitative results on HMDB51 [64], UCF101 [65](see §4.3). “INet-RN50” denotes ResNet-50 pre-trained on ImageNet. The best results are highlighted.
  HMDB51 UCF101
Method Reference Pre-training 1-shot 5-shot 1-shot 5-shot
AmeFuNet [17] MM20 INet-RN50 60.2 75.5 85.1 95.5
TRX [19] CVPR21 INet-RN50 - 75.6 - 96.1
TA2[66] AAAI22 INet-RN50 59.7 73.9 81.9 95.1
MTFAN [20] CVPR22 INet-RN50 59.0 74.6 84.8 95.1
HyRSM [13] CVPR22 INet-RN50 60.3 76.0 83.9 94.7
STRM [14] CVPR22 INet-RN50 52.3 77.3 80.5 96.9
HCL [21] ECCV22 INet-RN50 59.1 76.3 82.6 94.5
SloshNet [16] AAAI23 INet-RN50 - 77.5 - 97.1
MoLo [18] CVPR23 INet-RN50 60.8 77.4 86.0 95.5
GgHM [15] ICCV23 INet-RN50 85.2 96.3 61.2 76.9
CLIP-FSAR [61] IJCV24 CLIP-RN50 69.2 80.3 91.3 97.0
A2M2-Net [12] IJCV25 INet-RN50 61.8 76.6 86.1 95.8
CLIP-Freeze [57] ICML21 CLIP-ViT-B 58.2 77.0 89.7 95.7
CapFSAR [67] ArXiv23 CLIP-ViT-B 65.2 78.6 93.3 97.8
MVP-shot [68] TMM25 CLIP-ViT-B 77.0 88.1 96.8 99.0
MA-FSAR [58] PR25 CLIP-ViT-B 83.4 87.9 97.2 99.2
CLIP-FSAR [61] IJCV24 CLIP-ViT-B 77.1 87.7 97.0 99.1
TEAM [69] CVPR25 INet-ViT-B 70.9 85.5 94.5 98.8
HR2G-shot (ours) - CLIP-ViT-B $\mathbf{85.6}$ $\mathbf{88.6}$ $\mathbf{98.0}$ $\mathbf{99.3}$
4 Experiments
4.1 Experimental Setup
Dataset. We evaluate our HR2G-shot on five commonly used datasets, i.e., Kinetics [70], SSv2-small [71], SSv2-full [71], HMDB51 [64], and UCF101 [65]. For HMDB51 and UCF101, we adopt the few-shot split from [72, 73], with 3131/1010/$10$ classes and 7070/1010/2121 classes for train/val/test, respectively. Following previous works [18, 74, 75], we divide Kinetics, SSv2-small and SSv2-full into three splits: 6464/1212/$24$ classes used for train, val, and test, respectively. The difference between SSv2-small and SSv2-full is that SSv2-full contains more videos per class in train set.
Evaluation. Following standard evaluation protocols [10, 15, 18, 76, 77], 55-way 11-shot and $5$-shot accuracy over $10,000$ tasks are used for evaluation.
4.2 Implementation Details
Network Architecture. HR2G-shot adopts pre-trained CLIP ViT-B [57] as our backbone for parameter-efficient fine-tuning, for a fair comparison with previous methods [68, 61, 58]. In IKT, we empirically set the size of the temporal knowledge bank G=50G\!=\!50 (see Fig. 4). By default, the number of temporal prototypes $M$ is set to $3$ (see Table 8). For other hyper-parameters, we empirically set retrieval ratio $\kappa$ and momentum $\mu$ to $0.7$ (see Table 7) and $0.99$ (see Table 9), respectively.
Network Training. In line with previous works [10, 78, 61, 15], we uniformly and sparsely sample T=8T\!=\!8 frames from each video to encode frame-level representations, resizing each frame to a height of 256 pixels. During the training phase, we apply standard data augmentation techniques, including random horizontal flipping, cropping, and color jittering. Conversely, only center cropping is utilized during the testing phase to ensure deterministic evaluation. We freeze both the CLIP visual encoder and text encoder, and only finetune lightweight adapters in the visual encoder as [58]. The model is optimized using the Adam [79] optimizer, coupled with a multi-step learning rate scheduler. The entire training process spans $10$ epochs.
Table 2: Quantitative results on Kinetics [70] and SSv2-full [71] (see §4.3). “INet-RN50” denotes ResNet-50 pre-trained on ImageNet. The best results are highlighted.
  Kinetics SSv2-full
Method Reference Pre-training 1-shot 5-shot 1-shot 5-shot
CMN [9] ECCV18 INet-RN50 $60.5$ $78.9$ $36.2$ $48.9$
OTAM [10] CVPR20 INet-RN50 $73.0$ $85.8$ $42.8$ $52.3$
AmeFuNet [17] MM20 INet-RN50 $74.1$ $86.8$ - -
TRX [19] CVPR21 INet-RN50 $63.6$ $85.9$ - $64.6$
TA2[66] AAAI22 INet-RN50 $72.8$ $85.8$ $47.6$ $61.0$
MTFAN [20] CVPR22 INet-RN50 $74.6$ $87.4$ $45.7$ $60.4$
HyRSM [13] CVPR22 INet-RN50 $73.7$ $86.1$ $54.3$ $69.0$
STRM [14] CVPR22 INet-RN50 $62.9$ $86.7$ $43.1$ $68.1$
HCL [21] ECCV22 INet-RN50 $73.7$ $85.8$ $47.3$ $64.9$
SloshNet [16] AAAI23 INet-RN50 - $87.0$ - -
MoLo [18] CVPR23 INet-RN50 $74.0$ $85.6$ $55.0$ $69.6$
GgHM [15] ICCV23 INet-RN50 $74.9$ $87.4$ $54.5$ $69.2$
CLIP-FSAR [61] IJCV24 CLIP-RN50 $87.6$ $91.9$ $58.1$ $62.8$
A2M2-Net [12] IJCV25 INet-RN50 74.5 87.5 56.7 74.1
CLIP-Freeze [57] ICML21 CLIP-ViT-B $78.9$ $91.9$ $30.0$ $42.4$
CapFSAR [67] Arxiv23 CLIP-ViT-B $84.9$ $93.1$ $51.9$ $68.2$
MVP-shot [68] TMM25 CLIP-ViT-B $91.0$ $95.1$ - -
MA-FSAR [58] PR25 CLIP-ViT-B $\mathbf{95.7}$ $96.0$ $63.3$ $72.3$
CLIP-FSAR [61] IJCV24 CLIP-ViT-B $94.8$ $95.4$ $62.1$ $72.1$
TEAM [69] CVPR25 INet-ViT-B $83.3$ $92.9$ - -
HR2G-shot (ours) - CLIP-ViT-B $95.2$ $\mathbf{96.4}$ $\mathbf{65.4}$ $\mathbf{74.8}$
Table 3: Quantitative results on SSv2-small [71] (see §4.3). “INet-RN50” denotes ResNet-50 pre-trained on ImageNet. The best results are highlighted.
  SSv2-small
Method Reference Pre-training 1-shot 5-shot
CMN [9] ECCV18 INet-RN50 $34.4$ $43.8$
TRX [19] CVPR21 INet-RN50 - $59.1$
HyRSM [13] CVPR22 INet-RN50 $40.6$ $56.1$
STRM [14] CVPR22 INet-RN50 $37.1$ $55.3$
HCL [21] ECCV22 INet-RN50 $38.9$ $55.4$
MoLo [18] CVPR23 INet-RN50 $41.9$ $56.2$
CLIP-FSAR [61] IJCV24 CLIP-RN50 $52.0$ $55.8$
A2M2-Net [12] IJCV25 INet-RN50 42.9 60.5
CLIP-Freeze [57] ICML21 CLIP-ViT-B $29.5$ $42.5$
CapFSAR [67] Arxiv23 CLIP-ViT-B $45.9$ $59.9$
MVP-shot [68] TMM25 CLIP-ViT-B $55.4$ $62.0$
MA-FSAR [58] PR25 CLIP-ViT-B $59.1$ $64.5$
CLIP-FSAR [61] IJCV24 CLIP-ViT-B $54.6$ $61.8$
TEAM [69] CVPR25 INet-ViT-B $47.2$ $63.1$
HR2G-shot (ours) - CLIP-ViT-B $\mathbf{60.2}$ $\mathbf{66.0}$